# Complex-valued in-medium potential between heavy impurities in ultracold atoms

### v1

@@ III.2 Polarons in superfluid gas
As a second example, we consider impurities in superfluids. Such a system is realized as impurities immersed in either a BEC of the Bose gas [14, 15, 16, 17, 18] or a pair-condensed Fermi superfluid including the unitary Fermi gas.
To clarify the long-range behavior of $V_{\mathop{\mathrm{Im}}}(r)$ in superfluid medium, we rely on an effective Hamiltonian of a superfluid phonon originally developed by Landau in describing the superfluid ${}^{4}$e [54, 55]. One advantage of this formulation is that it captures a universal long-distance behavior—guaranteed by the Nambu-Goldstone theorem [56, 57, 58]—without relying on the perturbative treatment of the interaction between medium particles. As a consequence, the following result covers not only the weakly-coupled Bose superfluid but also the strongly-coupled Fermi superfluid across BEC-BCS crossover, including the unitary regime, at the cost of the ability to describe the short-distance behavior around, e.g., a healing length. Our formulation is thus complementary to more familiar formulations for the Bose polarons such as the analysis based on the so-called Fröhlich model [59, 60], or the variational wave functions [19, 14, 61, 62].
In the superfluid phase, we have the number density and phonon field as low-energy collective degrees of freedom as indicated by the spontaneously broken $\text{U}(1)$ symmetry. Let $\hat{n}(\bm{x})$ and $\hat{\varphi}(\bm{x})$ be such quantum operators, which satisfy the canonical commutation relation
$[\hat{n}(\bm{x}),\,\hat{\varphi}(\bm{x}^{\prime})]=-\mathrm{i}\delta^{(3)}(\bm{x}-\bm{x}^{\prime}).$ (34)
We then find the effective Hamiltonian of the superfluid phonon interacting with the impurity as follows [54, 55]:
@@
where $m$ is the mass of medium particles and $\epsilon(\hat{n})$ denotes an internal energy as a function of the number density. 111 The effective Hamiltonian (35) is equivalent to the effective Lagrangian employed in [38], from which the real part of the in-medium potential was derived. On the equivalence between these, we refer to Ref. [63].
Considering the fluctuation on the top of the global equilibrium, we expand $\hat{n}(\bm{x})=\bar{n}+\delta\hat{n}(\bm{x})$ and expand $\epsilon(\hat{n})$ at the quadratic order. As a result, the effective Hamiltonian turns into
@@
where we defined the inverse charge susceptibility $\chi^{-1}=\epsilon^{\prime\prime}({\bar{n}})$ and omit the constant term $\epsilon({\bar{n}})$ and the higher-order term with more than three $\delta\hat{n}$. Since we are interested in the long-distance behavior, we can regard the interaction term appearing in the first line of Eq. (36) as a perturbation. We emphasize that this treatment follows from the derivative expansion [63], and does not require a small coupling constant among medium particles. Based on this, we evaluate the retarded Green’s function for the number density operator $\hat{n}(\bm{x})$.
Again, we employ the imaginary-time formalism, in which we have a set of the Mastubara Green’s functions given by
@@
We now perform a perturbative expansion with respect to the medium interaction term (38). In this expansion, the leading one-loop diagram that contributes to the in-medium potential is given in the left diagram of Fig. 3. Recalling the zero-frequency limit in Eqs. (17) and (18), it is sufficient to evaluate the low-frequency behavior of the retarded Green’s function. At vanishing frequencies, the propagation of the density fluctuation $\delta\hat{n}(\bm{x})$ is suppressed as $\Delta_{nn}(\bm{k},\mathrm{i}\omega^{B}_{n}=0)=\chi$. Thereby, the Green’s function of the number density fluctuation directly connected to the impurity lines on the left diagram of Fig. 3 can be replaced with $\Delta_{nn}(\bm{k},\mathrm{i}\omega_{n})\simeq\chi$ at low frequencies. As a result, the resulting Feynman diagram contributing to the in-medium potential is shown on the left of Fig. 3, which agrees with the diagram computed in Ref. [38] to evaluate the real part of the potential.
Having identified the relevant one-loop diagram for the in-medium potential, we now find the corresponding Matsubara Green’s function as
@@
where we also added the tree-level contribution as the first term. Here, $n_{B}(E)\equiv 1/(\mathrm{e}^{\beta E}-1)$ is the Bose distribution function. Performing the analytic continuation and the Fourier transformation, we find the retarded Green’s function
$G^{R}(\bm{r},\omega)$ $=\chi\delta^{(3)}(\bm{r})-\frac{1}{2m^{2}}\int_{\bm{k},\bm{q}}\mathrm{e}^{\mathrm{i}(\bm{k}-\bm{q})\cdot\bm{r}}\frac{(\bm{q}\cdot\bm{k})^{2}}{4E_{\bm{k}}E_{\bm{q}}}\Bigg{[}\big{[}1+n_{B}(E_{\bm{k}})+n_{B}(E_{\bm{q}})\big{]}\left(\frac{1}{\omega+\mathrm{i}0-E_{\bm{k}}-E_{\bm{q}}}-\frac{1}{\omega+\mathrm{i}0+E_{\bm{k}}+E_{\bm{q}}}\right)$
$\hskip 136.0pt+\big{[}n_{B}(E_{\bm{k}})-n_{B}(E_{\bm{q}})\big{]}\left(\frac{1}{\omega+\mathrm{i}0+E_{\bm{k}}-E_{\bm{q}}}-\frac{1}{\omega+\mathrm{i}0-E_{\bm{k}}+E_{\bm{q}}}\right)\Bigg{]}.$ (40)
The first term corresponds to the well-known Yukawa potential [31], which is effectively squashed into a contact interaction in terms of the effective theory describing distances longer than the healing length.
The real part of the in-medium potential obtained from Eq. (40) agrees with that computed in Ref. [38] (except for the contact interaction arising from the first term), while we identify the imaginary part as
@@
×$\quad\times[1+n_{B}(E_{\bm{k}})]n_{B}(E_{\bm{k}}).$ (41)
Performing the angular integration with $g\simeq 2\pi a_{\mathrm{IM}}/m$, we eventually obtain the following result:
@@
where we introduced a thermal phonon momentum scale $k_{T}\equiv T/c_{s}$ and the function $h(y)$
@@
×$\quad\times\left[\frac{3j_{1}(sy)^{2}}{(sy)^{2}}-\frac{2j_{1}(sy)j_{2}(sy)}{sy}+j_{2}(sy)^{2}\right],$ (43)
with the spherical Bessel functions $j_{n}(x)$. The details of the calculations are found in Appendix. A
Let us specify asymptotic behaviors of $V_{\mathop{\mathrm{Im}}}(r)$ at short and long distances compared to a thermal phonon length scale $k_{T}^{-1}=c_{s}/T$. Using the asymptotic behavior of the function $h(y)$ (see Appendix. A), we find that $V_{\mathop{\mathrm{Im}}}(r)$ behaves at short and long distances as
@@
Figure 4 demonstrates the imaginary part of the induced potential in superfluids for several temperatures. In contrast to the Fermi polaron case, there is no oscillatory behavior and the magnitude of the potential is more sensitive to the temperature. One also explicitly confirms the $r^{-2}$ scaling at long distances, as clearly shown in the right panel of Fig. 4.
Refer to caption (a)
Refer to caption (b)
Figure 4: The imaginary part of the induced potential for polarons in the superfluid. (a) at short distances and (b) long distances. Black, blue, and red curves show results for $T/(mc_{s}^{2})=0.6,0.8$, and $1.0$. In panel (b), The dashed curve shows the power-law decay $r^{-2}$, which matches the asymptotic behaviors of the induced potential.
@@ III.3 Universal long-range power-law decay
As we will show below, the power-law behavior of $V_{\mathop{\mathrm{Im}}}(r)$ at long distances originates from the common structure of the low-energy scattering between the impurity and the medium excitation. To clarify this origin, let us consider the imaginary potential in momentum space
$\widetilde{V}^{F}_{\mathop{\mathrm{Im}}}(\bm{k})$ $\propto-g^{2}\int_{\bm{q}}\delta(\xi_{\bm{q}+\bm{k}}-\xi_{\bm{q}})n_{F}(\xi_{\bm{q}})\big{[}1-n_{F}(\xi_{\bm{q}})\big{]},$ (46)
@@
It is crucial to notice that the above results allow a scattering interpretation: The imaginary part is expressed in terms of the scattering amplitude of the medium excitation with the static impurity as shown in Fig. 5  (see, e.g., Ref. [64] for the so-called cutting rule). The integrand is composed of three parts; (a) the delta function imposing the energy conservation of the on-shell medium excitation, (b) the (properly normalized) scattering cross section part,222 Note that $E_{\bm{q}}^{-2}$ appears for the superfluid case from the proper normalization with $\delta(E_{\bm{q}+\bm{k}}-E_{\bm{q}})$. and (c) thermal distributions for incoming and outgoing excitations.
Figure 5: Scattering interpretation of the imaginary part of the induced potential. Taking the imaginary part results in cutting the diagram with the dashed line in the left panel. As a result, the imaginary part is given by the right panel, which gives a cross section of the on-shell medium excitation (solid lines) scattered by the heavy static impurity (amputated bold lines).
Let us now investigate the low-$\bm{k}$ behaviors of these building blocks. We first note that thermal distributions are $\bm{k}$-independent, and thus, the low-momentum behavior is controlled by the delta function and cross section. At low momentum $\bm{k}\simeq\bm{0}$, one can approximate the delta functions as
@@
yielding the $k^{-1}$ factor ($\theta_{\bm{q}}$ is the angle between $\bm{q}$ and $\bm{k}$, which is fixed $\theta_{\bm{q}}=\pi/2$ by performing the $\bm{q}$-integration). As a result, if the remaining part—the low-energy scattering cross section—approaches to the non-vanishing constant, the above $k^{-1}$ factor governs the long-distance behavior, leading to $r^{-2}$ behavior of $V_{\mathop{\mathrm{Im}}}(r)$ in coordinate space.
@@
Therefore, when the impurity and the medium excitations can collide elastically, the imaginary potential shows a universal power-law decay, $V_{\mathop{\mathrm{Im}}}(r)\propto r^{-2}$, at long distances. Such a condition is expected to hold true not only for the Fermi and Bose polaron systems presented above, but rather for a wide variety of physical systems. In fact, the example in subatomic physics—the in-medium heavy-quark potential mentioned above—fits into this class. On the other hand, we also note that the elastic scattering condition may be violated for not so heavy impurity, where the recoil effects are relevant [65].
@@ IV.2 Spectral width of bipolaron at low temperature
The imaginary potential between the polarons affects the spectral width of the bound state of two polarons, i.e., bipolaron. The imaginary part of the in-medium potential is smaller than the real part at low temperatures because it is proportional to the temperature as in Eq. (18). At this time, we can essentially regard the relative wave function of the bipolaron $\Psi_{b}(\bm{r})$ created by the attractive $V_{\mathop{\mathrm{Re}}}(\bm{r})$, and perturbatively estimate its spectral width $-\frac{1}{2}\Gamma$ as (see Appendix B for a more formal discussion)
@@
Let us crudely evaluate the spectral width. Because $V_{\mathop{\mathrm{Im}}}(\bm{r})$ has its maximum absolute value at $\bm{r}=\bm{0}$, the second term in Eq. (50) achieves that maximum $V_{\mathop{\mathrm{Im}}}(\bm{0})$ when the two impurities get close. Using Eq. (21), the width at that time is found as $\frac{1}{2}\Gamma\simeq-4E^{(1)}_{\mathop{\mathrm{Im}}}$. This result can be understood physically as follows: The medium recognizes the two polarons as a single molecule with its coupling constant $2g$. The imaginary part of the self-energy of this single molecule is given by that of one polaron with $g^{2}$ replaced by $(2g)^{2}$, resulting in the width of one molecule being four times that of one polaron. In the opposite case where the two polarons are far apart, the second term in Eq. (50) is negligible compared to the first term, and thus the width is given by the sum of the contributions of the two independent polarons as $\frac{1}{2}\Gamma\simeq-2E^{(1)}_{\mathop{\mathrm{Im}}}$. These two extreme cases give crude upper and lower bounds for the width as $2|E^{(1)}_{\mathop{\mathrm{Im}}}|<\frac{1}{2}\Gamma<4|E^{(1)}_{\mathop{\mathrm{Im}}}|$, where the explicit form of $2E^{(1)}_{\mathop{\mathrm{Im}}}=V_{\mathop{\mathrm{Im}}}(\bm{0})$ is found in Eqs. (31) and (44). This inequality holds true as long as the imaginary part can be treated perturbatively.
@@ IV.3 Medium density fluctuation induced by a single impurity
×$\times A_{0}^{\mathrm{eff}}(\bm{x}^{\prime},t^{\prime}).$ (51)
@@ V Conclusion and outlook
In this paper, we have studied the complex-valued in-medium potential between two heavy impurities immersed in quantum environments, with a particular focus on the Bose and Fermi polarons in cold atoms. We provided a general definition of the in-medium potential (3)-(4) from the long-time behavior of the correlation function for two immobile impurities immersed in the medium. When the impurity-medium coupling is a weak $s$-wave contact coupling, the real and imaginary parts of the complex-valued potential are expressed by the low-frequency limit of the two-point retarded Green’s functions of medium number density fluctuations as given in Eqs. (16)-(18).
@@ Acknowledgements
We thank Pak Hang Chris Lau and Tilman Enss for fruitful discussions. This work is supported by Japan Society for the Promotion of Science (JSPS) KAKENHI Grant Number JP23H01174, and partially by RIKEN iTHEMS Program and YITP domestic molecule-style workshop at Kyoto University. Y.A. is supported by JSPS KAKENHI Grant Number JP18K13538. S.E. is supported by JSPS KAKENHI Grant Numbers JP21H00116 and JP22K03492. K.F. is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), Project-ID 273811115 (SFB1225ISOQUANT). M.H. is supported by JSPS KAKENHI Grant Number 22K20369.
Appendix A Derivation of Eqs. (41)-(45)

### v2

@@ III.2 Polarons in superfluid gas
As a second example, we consider impurities in superfluids. Such a system is realized as impurities immersed in a BEC of the Bose gas [35, 32, 61, 79, 51], or a pair-condensed Fermi superfluid including the unitary Fermi gas.
To clarify the long-range behavior of $V_{\mathop{\mathrm{Im}}}(r)$ in superfluid medium, we rely on an effective Hamiltonian of a superfluid phonon originally developed by Landau in describing the superfluid 4He [39, 47]. One advantage of this formulation is that it captures a universal long-distance behavior—guaranteed by the Nambu-Goldstone theorem [55, 28, 29]—without relying on the perturbative treatment of the interaction between medium particles. As a consequence, the following result covers not only the weakly-coupled Bose superfluid but also the strongly-coupled Fermi superfluid across BEC-BCS crossover, including the unitary regime, at the cost of the ability to describe the short-distance behavior around, e.g., a healing length. Our formulation is thus complementary to more familiar formulations for the Bose polarons such as the analysis based on the so-called Fröhlich model [25, 27], or the variational wave functions [15, 35, 80, 72].
In the superfluid phase, we have the number density and phonon field as low-energy collective degrees of freedom as indicated by the spontaneously broken $\text{U}(1)$ symmetry. Let $\hat{n}(\bm{x})$ and $\hat{\varphi}(\bm{x})$ be such quantum operators, which satisfy the canonical commutation relation $\hat{n}(\bm{x}),\,\hat{\varphi}(\bm{x}^{\prime})]=-\mathrm{i}\delta^{(3)}(\bm{x}-\bm{x}^{\prime})$. We then find the effective Hamiltonian of the superfluid phonon interacting with the impurity as follows [39, 47]:
@@
[…]
@@
where we defined the inverse charge susceptibility $\chi^{-1}=\epsilon^{\prime\prime}({\bar{n}})$ and omit the constant term $\epsilon({\bar{n}})$ and the higher-order term with more than three $\delta\hat{n}$. Since we are interested in the long-distance behavior, we can regard the interaction term appearing in the first line of Eq. (24) as a perturbation. We emphasize that this treatment follows from the derivative expansion [75], and does not require a small coupling constant among medium particles.
To evaluate the retarded Green’s function for the number density operator $\hat{n}(\bm{x})$, we employ again the imaginary-time formalism:
@@
We now perform a perturbative expansion with respect to the medium interaction term (26): the leading one-loop diagram that contributes to the in-medium potential is given in the left diagram of Fig. 3. Recalling Eqs. (8) and (9), we evaluate the low-frequency behavior of the retarded Green’s function. At vanishing frequencies, the propagation of the density fluctuation $\delta\hat{n}(\bm{x})$ is suppressed as $\Delta_{nn}(\bm{k},\mathrm{i}\omega^{B}_{n}=0)=\chi$. Thereby, the Green’s function of the number density directly connected to the impurity lines on the left diagram of Fig. 3 can be replaced with $\Delta_{nn}(\bm{k},\omega)\simeq\chi$ after analytic continuation to a low real frequency. The resulting diagram is shown on the right of Fig. 3, which agrees with the diagram computed in Ref. [26] to evaluate the real part of the potential.
We thus find the corresponding Green’s function as
@@
where we also added the tree-level contribution as the first term. Here, $n_{B}(E)\equiv 1/(\mathrm{e}^{\beta E}-1)$ is the Bose distribution. Performing the analytic continuation and the Fourier transformation, we find the retarded Green’s function $G^{R}(\bm{r},\omega)$, from which the in-medium potential is obtained. The real part agrees with that computed in Ref. [26] (except for the contact interaction arising from the first term), while we identify the imaginary part as
@@
$\quad\times[1+n_{B}(E_{\bm{k}})]n_{B}(E_{\bm{k}}).$ (28)
Performing the angular integration with $g\simeq 2\pi a_{\mathrm{IM}}/m$, we obtain the following result (see Appendix. C):
@@
where $k_{T}\equiv T/c_{s}$ is a thermal phonon momentum and
@@
$\quad\times\left[\frac{3j_{1}(sy)^{2}}{(sy)^{2}}-\frac{2j_{1}(sy)j_{2}(sy)}{sy}+j_{2}(sy)^{2}\right],$ (30)
with the spherical Bessel functions $j_{n}(x)$.
Using the asymptotic behavior of $h(y)$ (see Appendix. C), we find that $V_{\mathop{\mathrm{Im}}}(r)$ behaves at short and long distances (compared to a thermal phonon length scale $k_{T}^{-1}=c_{s}/T$) as
@@
Figure 4 shows $V_{\mathop{\mathrm{Im}}}(r)$ in superfluids for several temperatures. In contrast to the Fermi polaron case, there is no oscillatory behavior and the magnitude of the potential is more sensitive to the temperature. One also explicitly confirms the $r^{-2}$ scaling at long distances, as clearly shown in the right panel of Fig. 4.
Figure 4: The imaginary part of the induced potential for polarons in the superfluid. (a) at short distances and (b) long distances. Solid blue, dashed-dotted green, and dashed red curves show results for $T/(mc_{s}^{2})=0.6,0.8$, and $1.0$. In panel (b), the dotted black curve shows the power-law decay $r^{-2}$, which matches the asymptotic behaviors of the induced potential.
@@ III.3 Universal long-range power-law decay
The power-law behavior originates from the common structure of the low-energy scattering between the impurity and the medium excitation. To clarify this origin, let us consider the imaginary potential in momentum space
$\widetilde{V}^{F}_{\mathop{\mathrm{Im}}}(\bm{k})$ $\propto-g^{2}\int_{\bm{q}}\hskip-2.0pt\delta(\xi_{\bm{q}+\bm{k}}-\xi_{\bm{q}})n_{F}(\xi_{\bm{q}})\big[1-n_{F}(\xi_{\bm{q}})\big],$ (33)
@@
These equations allow a scattering interpretation: The imaginary part is expressed in terms of the scattering amplitude of the medium excitation with the static impurity as shown in Fig. 5  (see, e.g., Ref. [45] for the so-called cutting rule). The integrand is composed of three parts; (a) the delta function imposing the energy conservation of the on-shell medium excitation, (b) the (properly normalized) scattering cross section part, [^2] Note that $E_{\bm{q}}^{-2}$ appears for the superfluid case from the proper normalization with $\delta(E_{\bm{q}+\bm{k}}-E_{\bm{q}})$. and (c) thermal distributions for incoming and outgoing excitations.
Figure 5: Scattering interpretation of the imaginary part of the induced potential. Taking the imaginary part results in cutting the diagram with the dashed line in the left panel. The imaginary part is thus given by the right panel, which gives a cross section of the on-shell medium excitation (solid lines) scattered by the static impurity (amputated bold lines).
Let us now investigate the low-$\bm{k}$ behaviors of these building blocks. Because the thermal distributions are $\bm{k}$-independent, the low-momentum behavior is controlled by the delta function and cross section. At low momentum $\bm{k}\simeq\bm{0}$, one can approximate the delta functions as
@@
yielding the $k^{-1}$ factor ($\theta_{\bm{q}}$ is the angle between $\bm{q}$ and $\bm{k}$, which is fixed $\theta_{\bm{q}}=\pi/2$ by performing the $\bm{q}$-integration). As a result, if the low-energy scattering cross section approaches to the non-vanishing constant, the above $k^{-1}$ factor governs the long-distance behavior, leading to $r^{-2}$ behavior of $V_{\mathop{\mathrm{Im}}}(r)$ in coordinate space.
@@
Therefore, when the impurity and the medium excitations can collide elastically, the imaginary potential shows a universal power-law decay, $V_{\mathop{\mathrm{Im}}}(r)\propto r^{-2}$, at long distances. Such a condition is expected to hold true not only for the Fermi and Bose polaron systems presented above, but rather for a wide variety of physical systems. In fact, the example in subatomic physics—the in-medium heavy-quark potential mentioned above—fits into this class. On the other hand, the elastic scattering condition may be violated for not so heavy impurity, where the recoil effects are relevant [66].
@@ IV.2 Spectral width of bipolaron at low temperature
The imaginary potential between the polarons affects the spectral width of the bound state of two polarons, i.e., bipolaron. The imaginary part of the in-medium potential is smaller than the real part at low temperatures because it is proportional to the temperature as in Eq. (9). We can thus essentially regard the relative wave function of the bipolaron $\Psi_{b}(\bm{r})$ created by the attractive $V_{\mathop{\mathrm{Re}}}(\bm{r})$, and perturbatively estimate its spectral width $\frac{1}{2}\Gamma$ as (see Appendix B for a more formal discussion)
@@
[…]
@@ IV.3 Medium density fluctuation induced by a single impurity
$\times A_{0}^{\mathrm{eff}}(\bm{x}^{\prime},t^{\prime}).$ (38)
@@ V Conclusion and outlook
We have studied the complex-valued in-medium potential between two heavy impurities immersed in quantum environments, with a particular focus on the Bose and Fermi polarons in cold atoms. We provided a general definition of the in-medium potential (3)-(4) from the long-time behavior of the correlation function for two immobile impurities immersed in the medium. When the impurity-medium coupling is a weak $s$-wave contact coupling, the real and imaginary parts of the complex-valued potential are expressed by the low-frequency limit of the two-point retarded Green’s functions of medium number density fluctuations as given in Eqs. (9)-(10).
@@ Acknowledgements
We thank Pak Hang Chris Lau and Tilman Enss for fruitful discussions. This work is supported by Japan Society for the Promotion of Science (JSPS) KAKENHI Grant Number JP23H01174, and partially by RIKEN iTHEMS Program and YITP domestic molecule-style workshop at Kyoto University. Y.A. is supported by JSPS KAKENHI Grant Number JP18K13538. S.E. is supported by JSPS KAKENHI Grant Numbers JP21H00116 and JP22K03492. K.F. is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), Project-ID 273811115 (SFB1225ISOQUANT) and JSPS KAKENHI Grant Number 24KJ0062. M.H. is supported by JSPS KAKENHI Grant Number 22K20369.
Appendix A Derivation of the formula (8)-(12) in the weak impurity-medium coupling case
We here provide a detailed derivation of the perturbative formula (8)-(12) for the complex-valued potential for the interested readers. The derivation in this appendix serves as an altenative proof of those formulas based on the operator formalism rather than the path-integral formalism in Ref. [73].
For a system with weak impurity-medium coupling, defined by Eqs. (6) and (7) in the main text, the interaction picture with the choice of the unperturbed Hamiltonian
with $\mathrm{T}$ denoting the time-ordered product. Here, we introduced $\hat{\mathcal{O}}(t)\equiv\mathrm{e}^{\mathrm{i}t\hat{H}_{0}}\hat{\mathcal{O}}\mathrm{e}^{-\mathrm{i}t\hat{H}_{0}}$ for an arbitrary operator $\hat{\mathcal{O}}$, and the number density fluctuation operator around its thermal average as $\delta\hat{n}(\bm{x})=\hat{n}(\bm{x})-\langle\hat{n}(\bm{x})\rangle$.
We take the heavy-impurity limit as the first step to evaluate $\Psi(\bm{r},t)$. In this limit, the large impurity mass $M_{*}^{-1}\simeq M^{-1}$ suppresses its kinetic terms, i.e. the impurity is immobile and can be treated as a test particle fixed at a certain position. Then, the impurity density operator $\hat{\Phi}^{\dagger}(\bm{x},t)\hat{\Phi}(\bm{x},t)$ is time-independent and we obtain
To proceed further, we employ the weak-coupling expansion with respect to the impurity-medium coupling $g$. In the ladder approximation, we get
where the translational and rotational invariance of the medium is used. Matching this expression with the aforementioned long-time behavior $\Psi(\bm{r},t)\simeq\mathrm{e}^{-\mathrm{i}\bar{V}(\bm{r})t}\Psi(\bm{r},0)$, we find $\bar{V}(\bm{r})$ at $O(g^{2})$ as
According to Eq. (4), subtracting the one-body energy, we arrive at the expression of the in-medium interaction potential $V(\bm{r})$ for a weak impurity-medium coupling case as
$V(\bm{r})$ $\simeq\lim_{t\to\infty}\frac{g^{2}}{\mathrm{i}t}\int_{0}^{t}\mathrm{d}t_{1}\hskip-2.0pt\int_{0}^{t}\mathrm{d}t_{2}\,\langle\mathrm{T}\delta\hat{n}(\bm{r},t_{1})\delta\hat{n}(\bm{0},t_{2})\rangle,$
where we dropped $O(g^{3})$ terms and assumed that $\langle\delta\hat{n}(\bm{r},s)\delta\hat{n}(\bm{0},0)\rangle$ at fixed $\bm{r}$ decays in a finite duration of time. We here note that the zero temperature limit of Eq. (47), expressed in terms of the time-ordered product, reproduces a familiar formula of the induced potential at weak coupling [23].
By separating the real and imaginary parts of Eq. (47), we can express them with the retarded Green’s function. For instance, recalling $\langle\hat{A}\hat{B}\rangle^{*}=\langle\hat{B}^{\dagger}\hat{A}^{\dagger}\rangle$ and using the fluctuation-dissipation relation, we find
$V_{\mathrm{Im}}(\bm{r})$ $=-\frac{g^{2}}{2}\hskip-3.0pt\int_{-\infty}^{\infty}\hskip-7.0pt\mathrm{d}s\,\Big\langle\delta\hat{n}(\bm{r},s)\delta\hat{n}(\bm{0},0)+\delta\hat{n}(\bm{0},0)\delta\hat{n}(\bm{r},s)\Big\rangle=-g^{2}\lim_{\omega\to 0}\frac{2T}{\omega}\mathop{\mathrm{Im}}G^{R}(\bm{r},\omega).$ (48)
This completes the derivation of the perturbative formulas (8)-(12) in the main text.
Appendix B Complex potential in open systems
where the kernel $\mathcal{K}$ is given by
where we use our notation for the medium correlation functions. The higher order terms in the derivative expansion are omitted for simplicity. From this Markovian influence functional, one can derive quantum master equation for $N$ impurities. In Ref. [73], the authors take a classical limit to obtain Fokker-Planck equation, from which they further get generalized Langevin equation for the impurities. For details, we refer to [8] as well as [73], latter of which uses the formalism developed in the former.
where $\theta(\bm{x},t)$ is a Gaussian white noise with finite correlation length [note that $V_{\mathop{\mathrm{Im}}}(\bm{x})$ is positive definite] satisfying
$\bm{f}(\bm{r}_{i},t)=-\bm{\nabla}\theta(\bm{r}_{i},t),$ (55)
the correlation of random force is
After some algebra, we can get the transition rate
where $\bm{r}_{\pm}=\frac{1}{2}(\bm{r}\pm\bm{r}^{\prime})$ and $\Gamma$ is given in Eq. (37). The second term subtracts from the total width $\Gamma$ the rate of staying in $\Psi_{b}$ after a collision, which approaches $\Gamma$ when the bound state size is small. Therefore, the transition rate is actually smaller than the spectral width. Similar calculation was done in quarkonium in the quark-gluon plasma [36], but the total width and the transition rate are equal (i.e., the second term vanishes) because of the opposite charges of the heavy quark-antiquark pair.
Appendix C Derivation of Eqs. (29)-(32)